# A Delayed Radio Flare Traces Kinetic Energy Injection in the SMBHB Candidate SDSS J143016.05+230344.4

### 07 Mar 2026 v1

@@ Appendix A VLBI Observations and data reduction
Raw data were correlated with DiFX correlator (Deller et al., 2011) at the VLBA Data Correlation Center in Socorro, USA, with an integration time of 2 s and with the SFXC at Joint Institute for VLBI European Research Infrastructure Consortium (JIVE) in Dwingeloo, the Netherlands for the EVN (integration time 1 s). The correlated visibilities were then transferred to the China SKA Regional Centre (An et al., 2019) and processed with our automated VLBI pipeline [^1] https://github.com/SHAO-SKA/vlbi-pipeline, a Python+Parseltongue interface to AIPS (Greisen, 2003) following the standard AIPS Cookbook  [^2] http://www.aips.nrao.edu/cook.html.
@@ Appendix B MCMC fitting of VLBI core parameters
$\mathcal{V}_{\mathrm{m}}=f_{\nu}e^{-2\pi^{2}\left(\frac{\theta}{\sqrt{8\ln 2}}\right)^{2}\left(u^{2}+v^{2}\right)-2\pi i(u\rho+v\delta)}$ (B1)
@@
$\begin{array}[]{ll}\ln\mathcal{L}(x)=&-\frac{1}{2}\sum_{i=0}^{N}w_{i}\left[\left(\mathcal{V}_{R,\mathrm{m}}(u_{i},v_{i},x)-\mathcal{V}_{R,i}\right)^{2}\right.\\ &+\left.\left(\mathcal{V}_{I,\mathrm{m}}(u_{i},v_{i},x)-\mathcal{V}_{I,i}\right)^{2}\right]\end{array}$ (B2)
@@
We sample the posterior with emcee (Foreman-Mackey et al., 2013) using 8–16 walkers and $10^{4}$ steps (discarding the first half as burn-in). Priors are uniform and conservative: $0<f_{\nu}<5$ mJy, $0<\theta<2$ mas, and $|\rho|,|\delta|<2$ mas. Chains are initialized near the Difmap point-source solution with small Gaussian jitter. Convergence is verified by integrated autocorrelation times and by consistency of split-chain posteriors.
@@ C.1 Brightness Temperature
$T_{\rm B}=1.22\times 10^{12}(1+z)\left(\frac{S_{\nu}}{\rm Jy}\right)\left(\frac{\nu}{\rm GHz}\right)^{-2}\left(\frac{\theta}{\rm mas}\right)^{-2}{\rm~K},$ (C1)
@@
Because the core remains strictly unresolved across the high-frequency VLBA observations, the true emission region is intrinsically smaller than the synthesized beam. Consequently, the MCMC-derived structural upper limits on $\theta$ dictate strict lower limits on the brightness temperature ($T_{\rm B}\propto\theta^{-2}$). For instance, applying the rigorous MCMC 95% upper limit for the 15.2 GHz detection in early 2024 ($S_{\nu}\approx 0.305$ mJy, $\theta<0.22$ mas), Equation C1 implies $T_{\rm B}\gtrsim 3.6\times 10^{7}$ K. At lower frequencies where the size limits are more constraining relative to the flux (e.g., at 5 GHz), these lower limits securely exceed $\sim 10^{8}$ K. These values are highly characteristic of a compact, non-thermal AGN core and strongly disfavors thermal origins (which are restricted to $T_{\rm B}\lesssim 10^{5}$ K).
@@ C.2 Steady+flare SSA fitting
$S(\nu)=S_{p}\left(\frac{\nu}{\nu_{p}}\right)^{5/2}\left[1-\exp\!\left(-\tau(\nu)\right)\right],\qquad\tau(\nu)=\left(\frac{\nu}{\nu_{p}}\right)^{-(p+4)/2},$ (C2)
@@ C.3 Turnover-based equipartition scalings
For a homogeneous SSA component with electron index $p\simeq 3$, the equipartition radius and magnetic field can be robustly estimated from the turnover frequency $\nu_{p}$ and peak flux density $S_{p}$ (e.g., Readhead, 1994; Chevalier, 1998). Using the convenient normalizations in Chevalier (1998),
$R_{\rm eq}\simeq 4.0\times 10^{14}\left(\frac{f}{0.5}\right)^{-1/19}\left(\frac{S_{p}}{\rm Jy}\right)^{9/19}\left(\frac{D_{L}}{\rm Mpc}\right)^{18/19}\left(\frac{\nu_{p}}{5\rm~GHz}\right)^{-1}(1+z)^{-1}{\rm~cm},$ (C3)
$B_{\rm eq}\simeq 1.1\left(\frac{f}{0.5}\right)^{-4/19}\left(\frac{S_{p}}{\rm Jy}\right)^{-2/19}\left(\frac{D_{L}}{\rm Mpc}\right)^{-4/19}\left(\frac{\nu_{p}}{5\rm~GHz}\right)(1+z){\rm~G}.$ (C4)
@@ C.4 Cooling and expansion timescales
$t_{\rm syn}\approx 580{\rm~d~}\left(\frac{B}{1\rm~G}\right)^{-3/2}\left(\frac{\nu}{1\rm~GHz}\right)^{-1/2}(1+z)^{1/2}.$ (C5)
At the peak of the high-frequency outburst (Epoch II), our equipartition estimate yields $B_{\rm eq}\approx 1.35$ G. This implies a characteristic synchrotron cooling time of $t_{\rm syn}(15{\rm~GHz})\approx 1.0\times 10^{2}$ days. In contrast, the adiabatic expansion timescale is $t_{\rm ad}\sim R_{\rm eq}/v\approx 3$ days, assuming an initial outflow velocity of $v=0.2c$ at $R_{\rm eq}=4.53\times 10^{-4}$ pc.
@@ C.5 Apparent-speed limits from VLBI
$\beta_{\rm app}=\frac{\Delta\theta\,(1.526\ {\rm pc\,mas^{-1}})}{\Delta t_{\rm src}\,(0.3066\ {\rm pc\,yr^{-1}}c^{-1})}.$ (C6)
@@
$\Delta\theta_{\rm lim}\sim\max\!\left(0.3\text{--}0.5\,\theta_{\rm beam,min},\ 3\sigma_{\rm pos}\right),\quad\sigma_{\rm pos}\approx\frac{\theta_{\rm beam}}{2\,{\rm S/N}}.$ (C7)
Given our 15.2 GHz restoring beam minor axis of $\theta_{\rm beam,min}\sim 0.5$ mas and typical signal-to-noise ratios of S/N $\sim 5$99, the minimum resolvable separation limit falls in the range of $\Delta\theta_{\rm lim}\approx 0.15$$0.25$ mas. Substituting this angular limit into Equation C6 yields an apparent speed upper limit of $\beta_{\rm app}\lesssim 0.47$$0.79$, conservatively capping the bulk proper motion at $\lesssim 0.5$0.8c0.8\,c.
@@ Appendix D CNM density profile from the radio spectrum
$n\simeq\frac{B^{2}}{9\pi\,\epsilon_{B}\,m_{p}\,v^{2}}.$ (D1)

### 19 Mar 2026 v2

@@ Appendix A VLBI Observations and data reduction
Raw data were correlated with DiFX correlator (Deller et al. 2011) at the VLBA Data Correlation Center in Socorro, USA, with an integration time of 2 s and with the SFXC at Joint Institute for VLBI European Research Infrastructure Consortium (JIVE) in Dwingeloo, the Netherlands for the EVN (integration time 1 s). The correlated visibilities were then transferred to the China SKA Regional Centre (An et al. 2019) and processed with our automated VLBI pipeline [^1] https://github.com/SHAO-SKA/vlbi-pipeline, a Python+Parseltongue interface to AIPS (Greisen 2003) following the standard AIPS Cookbook  [^2] http://www.aips.nrao.edu/cook.html.
@@ Appendix B MCMC fitting of VLBI core parameters
$\mathcal{V}_{\mathrm{m}}=f_{\nu}e^{-2\pi^{2}\left(\frac{\theta}{\sqrt{8\ln 2}}\right)^{2}\left(u^{2}+v^{2}\right)-2\pi i(u\rho+v\delta)}$ (8)
@@
$\begin{array}[]{ll}\ln\mathcal{L}(x)=&-\frac{1}{2}\sum_{i=0}^{N}w_{i}\left[\left(\mathcal{V}_{R,\mathrm{m}}(u_{i},v_{i},x)-\mathcal{V}_{R,i}\right)^{2}\right.\\ &+\left.\left(\mathcal{V}_{I,\mathrm{m}}(u_{i},v_{i},x)-\mathcal{V}_{I,i}\right)^{2}\right]\end{array}$ (9)
@@
We sample the posterior with emcee (Foreman-Mackey et al. 2013) using 8–16 walkers and $10^{4}$ steps (discarding the first half as burn-in). Priors are uniform and conservative: $0<f_{\nu}<5$ mJy, $0<\theta<2$ mas, and $|\rho|,|\delta|<2$ mas. Chains are initialized near the Difmap point-source solution with small Gaussian jitter. Convergence is verified by integrated autocorrelation times and by consistency of split-chain posteriors.
@@ C.1 Brightness Temperature
$T_{\rm B}=1.22\times 10^{12}(1+z)\left(\frac{S_{\nu}}{\rm Jy}\right)\left(\frac{\nu}{\rm GHz}\right)^{-2}\left(\frac{\theta}{\rm mas}\right)^{-2}{\rm\penalty 10000\ K},$ (10)
@@
Because the core remains strictly unresolved across the high-frequency VLBA observations, the true emission region is intrinsically smaller than the synthesized beam. Consequently, the MCMC-derived structural upper limits on $\theta$ dictate strict lower limits on the brightness temperature ($T_{\rm B}\propto\theta^{-2}$). For instance, applying the rigorous MCMC 95% upper limit for the 15.2 GHz detection in early 2024 ($S_{\nu}\approx 0.305$ mJy, $\theta<0.22$ mas), Equation 10 implies $T_{\rm B}\gtrsim 3.6\times 10^{7}$ K. At lower frequencies where the size limits are more constraining relative to the flux (e.g., at 5 GHz), these lower limits securely exceed $\sim 10^{8}$ K. These values are highly characteristic of a compact, non-thermal AGN core and strongly disfavors thermal origins (which are restricted to $T_{\rm B}\lesssim 10^{5}$ K).
@@ C.2 Steady+flare SSA fitting
$S(\nu)=S_{p}\left(\frac{\nu}{\nu_{p}}\right)^{5/2}\left[1-\exp\!\left(-\tau(\nu)\right)\right],\qquad\tau(\nu)=\left(\frac{\nu}{\nu_{p}}\right)^{-(p+4)/2},$ (11)
@@ C.3 Turnover-based equipartition scalings
For a homogeneous SSA component with electron index $p\simeq 3$, the equipartition radius and magnetic field can be robustly estimated from the turnover frequency $\nu_{p}$ and peak flux density $S_{p}$ (e.g., Readhead 1994; Chevalier 1998). Using the convenient normalizations in Chevalier (1998),
$R_{\rm eq}\simeq 4.0\times 10^{14}\left(\frac{f}{0.5}\right)^{-1/19}\left(\frac{S_{p}}{\rm Jy}\right)^{9/19}\left(\frac{D_{L}}{\rm Mpc}\right)^{18/19}\left(\frac{\nu_{p}}{5\rm\penalty 10000\ GHz}\right)^{-1}(1+z)^{-1}{\rm\penalty 10000\ cm},$ (12)
$B_{\rm eq}\simeq 1.1\left(\frac{f}{0.5}\right)^{-4/19}\left(\frac{S_{p}}{\rm Jy}\right)^{-2/19}\left(\frac{D_{L}}{\rm Mpc}\right)^{-4/19}\left(\frac{\nu_{p}}{5\rm\penalty 10000\ GHz}\right)(1+z){\rm\penalty 10000\ G}.$ (13)
@@ C.4 Cooling and expansion timescales
$t_{\rm syn}\approx 580{\rm\penalty 10000\ d\penalty 10000\ }\left(\frac{B}{1\rm\penalty 10000\ G}\right)^{-3/2}\left(\frac{\nu}{1\rm\penalty 10000\ GHz}\right)^{-1/2}(1+z)^{1/2}.$ (14)
At the peak of the high-frequency outburst (Epoch II), our equipartition estimate yields $B_{\rm eq}\approx 1.35$ G. This implies a characteristic synchrotron cooling time of $t_{\rm syn}(15{\rm\penalty 10000\ GHz})\approx 1.0\times 10^{2}$ days. In contrast, the adiabatic expansion timescale is $t_{\rm ad}\sim R_{\rm eq}/v\approx 3$ days, assuming an initial outflow velocity of $v=0.2c$ at $R_{\rm eq}=4.53\times 10^{-4}$ pc.
@@ C.5 Apparent-speed limits from VLBI
$\beta_{\rm app}=\frac{\Delta\theta\,(1.526\ {\rm pc\,mas^{-1}})}{\Delta t_{\rm src}\,(0.3066\ {\rm pc\,yr^{-1}}c^{-1})}.$ (15)
@@
$\Delta\theta_{\rm lim}\sim\max\!\left(0.3\text{--}0.5\,\theta_{\rm beam,min},\ 3\sigma_{\rm pos}\right),\quad\sigma_{\rm pos}\approx\frac{\theta_{\rm beam}}{2\,{\rm S/N}}.$ (16)
Given our 15.2 GHz restoring beam minor axis of $\theta_{\rm beam,min}\sim 0.5$ mas and typical signal-to-noise ratios of S/N $\sim 5$99, the minimum resolvable separation limit falls in the range of $\Delta\theta_{\rm lim}\approx 0.15$$0.25$ mas. Substituting this angular limit into Equation 15 yields an apparent speed upper limit of $\beta_{\rm app}\lesssim 0.47$$0.79$, conservatively capping the bulk proper motion at $\lesssim 0.5$0.8c0.8\,c.
@@ Appendix D CNM density profile from the radio spectrum
$n\simeq\frac{B^{2}}{9\pi\,\epsilon_{B}\,m_{p}\,v^{2}}.$ (17)